# Governing Dynamic Capabilities: Cryptographic Binding and Reproducibility Verification for AI Agent Tool Use

### 15 Mar 2026 v1

@@ 4.3 Reproducibility as a Trust Primitive
The threshold $\theta$ is task-type dependent: our empirical validation (Experiment 5, §7) shows that code-generating agents achieve perfect determinism ($\theta=0.98$), while prose-generating agents require $\theta=0.10$ to account for paraphrase-level variation. Cross-model divergence remains well below even the relaxed threshold (max 0.089).
@@ 4.6 Security Analysis
Theorem 1 (Reproducibility Soundness, informal).
@@
By the DiFR token divergence result [1], the match rate between different models is $\delta\approx 0.06$ (character-level, validated in Experiment 5). By Hoeffding’s inequality: $\Pr[\mathrm{CharMatch}\geq\theta]\leq\exp(-2n(\theta-\delta)^{2})$, which is negligible for realistic output lengths.
@@ 4.7 Indistinguishability Under Sufficient Verification
$\epsilon\leq\exp\!\left(-2n\theta^{2}\right)$
Equivalently, for target confidence $1-\epsilon$, $n\geq\frac{\ln(1/\epsilon)}{2\theta^{2}}$ verification prompts suffice.
Proof sketch. Let $Z_{i}=\mathbf{1}[\mathrm{CharMatch}(M(x_{i}),M^{\prime}(x_{i}))<\theta]$ be independent Bernoulli random variables with $p=\Pr_{x\sim\mathcal{D}}[\mathrm{CharMatch}<\theta]$. Observing $\sum Z_{i}=0$ (all $n$ tests pass), by a one-sided binomial test: $\Pr[\sum Z_{i}=0\mid p\geq\epsilon]=(1-p)^{n}\leq(1-\epsilon)^{n}\leq\exp(-n\epsilon)$. Setting this equal to $\alpha$ (significance level) and solving yields $p\leq\ln(1/\alpha)/n$ with confidence $1-\alpha$. The Hoeffding refinement accounting for the threshold gap gives the stated bound. Full proof in §6. ∎
Practical implications. For code-generating agents ($\theta=0.98$), only $n=3$ prompts per skill achieve $\epsilon<10^{-2}$; $n=12$ achieves $\epsilon<10^{-10}$. Even for relaxed prose thresholds ($\theta=0.10$), $n=1200$ prompts achieve $\epsilon<10^{-10}$. These are operationally feasible: a verification suite of 50 prompts per skill, run periodically, provides strong guarantees.
Security interpretation. This transforms reproducibility from a detection mechanism into a safety certificate: an adversary who fine-tunes $M$ into $M^{\prime}$ that passes $n$ diverse replay checks has produced a model that is provably close to $M$ on the tested distribution. Any residual divergence is bounded by $\epsilon$, which decreases exponentially with the verification budget. The adversary faces a dilemma: either $M^{\prime}$ behaves like $M$ (and the fine-tuning achieves nothing), or $M^{\prime}$ diverges on some inputs and will be caught with probability increasing exponentially in $n$. This is analogous to cryptographic indistinguishability: the adversary’s advantage is negligible given sufficient verification.
@@ 4.8 Chain Verifiability
The gap that G2 uniquely fills. Consider what G1+G3 alone guarantee. G1 (capability binding) ensures that each agent’s certificate is bound to its declared model and tool set; any capability change invalidates the certificate. G3 (interaction auditability) ensures that every input and output is bilaterally signed and hash-chain linked; no record can be silently modified. Together, G1+G3 provide attribution integrity: we know who sent what to whom, and that no message was altered in transit.
@@
Principle 1: High-determinism models at interior nodes. Models with strong reproducibility (Anthropic Haiku: CharMatch==0.987, Sonnet: 0.823; see Table 4) should be placed at coordinator and interior positions in the trust tree, where their outputs become inputs for downstream agents. Low-determinism models (Kimi: 0.169, DeepSeek: 0.225) should be constrained to leaf positions where their outputs do not propagate further. This is not merely a performance recommendation—it is a security architecture constraint.
@@
Principle 2: Verification boundary checkpoints. When a G2==none agent must appear at an interior position (e.g., a reasoning model as coordinator), the chain requires a verification boundary: either (a) a human-in-the-loop checkpoint that re-anchors trust by reviewing and signing the output, (b) TEE attestation providing hardware-enforced computational integrity, or (c) dual-execution with independent instances and output comparison (analogous to Byzantine fault tolerance). Each mechanism re-establishes chain verifiability for downstream nodes.
@@
Principle 3: Explicit degradation in the ledger. When chain verifiability is broken at node $v_{k}$, the ledger should record a PARTIAL_VERIFIABILITY marker at that boundary. Downstream agents inherit a trust tier penalty beyond the standard G2==none downgrade (§4.4), making the reduced security guarantee visible to auditors and compliance systems.
Relationship to TEE. This analysis clarifies the complementary roles of G2 and trusted execution environments. G2 provides software-only verification of computational fidelity—a “software TEE”—using only black-box API access and the near-determinism of LLM inference. Hardware TEEs provide the same guarantee through hardware attestation but require infrastructure unavailable in most LLM deployments. The Chain Verifiability Theorem shows precisely where each mechanism is needed: at every interior agent node in the trust tree. In mixed deployments, G2 and TEE are interchangeable at the chain level: an agent with TEE attestation and G2==none maintains chain verifiability identically to an agent with G2==full and no TEE.
@@ 5.1 Motivation
Why OpenTelemetry traces are insufficient. The industry standard for agentic observability is OpenTelemetry (OTEL) [46], whose span trees record execution traces with span_id/parent_id linkage, timestamped input.value/output.value payloads, and span-kind annotations (CHAIN, LLM, TOOL). Analysis of the Nemotron-AIQ Agentic Safety Dataset [45]—10,796 OTEL traces from NVIDIA’s AI-Q research assistant under 22 distinct attack templates—reveals three critical gaps: (1) No integrity protection: span linkage is purely referential; any party with storage access can modify, delete, or reorder spans undetectably. (2) No identity binding: spans record function names but not cryptographically verified agent identities or capability certificates. (3) No privacy preservation: OTEL stores full plaintext input/output payloads, creating data exposure risk. Our interaction ledger addresses all three gaps through hash-chain linking ($h_{\mathit{prev}}$), Ed25519 bilateral signing ($\mathit{sig}_{i}$), agent certificate binding ($\mathit{ch}_{\mathit{send}},\mathit{ch}_{\mathit{recv}}$), and commitment-only storage ($\eta_{\mathit{in}},\eta_{\mathit{out}}$).
@@ 5.4 Forensic Reconstruction Protocol
6 Formal Security Properties and Proofs
We prove nine properties under the adversary model of §2, using standard cryptographic assumptions: EUF-CMA security, collision resistance of SHA-256, and the DiFR token divergence property [1].
Theorem 4 (Trust Containment).
@@ 6 Formal Security Properties and Proofs
Proof.
Theorem 5 (Capability Binding).
@@
Proof.
By collision resistance: $\Pr[H_{v}(t^{\prime})=H_{v}^{\mathit{cert}}]\leq\mathrm{negl}(\lambda)$ when $S_{v}(t^{\prime})\neq\sigma_{v}$. Forging a new certificate requires the parent’s key (same argument as Theorem 4). ∎∎
Theorem 6 (Chain Integrity).
@@
Proof.
Theorem 7 (Delegation Depth Enforcement).
@@
Proof.
Theorem 8 (Reproducibility Soundness).
If $v$ executes $M^{\prime}\neq M$ or $S^{\prime}\neq S$, replay verification detects with probability $\geq 1-\epsilon(\lambda)$.
Proof.
By DiFR [1]: $\mathbb{E}[\mathrm{CharMatch}(O^{\mathit{actual}},O^{\mathit{ref}})]\leq\delta$ for different models. By Hoeffding: $\Pr[\mathrm{CharMatch}\geq\theta]\leq\exp(-2n(\theta-\delta)^{2})$, negligible for realistic $n$. ∎∎
@@
If $M^{\prime}$ passes replay verification on $n$ independent prompts from $\mathcal{D}$, then $\Pr_{x\sim\mathcal{D}}[\mathrm{CharMatch}(M(x),M^{\prime}(x))<\theta]\leq\epsilon$ where $\epsilon\leq\exp(-2n\theta^{2})$.
Proof.
Let $p=\Pr_{x\sim\mathcal{D}}[\mathrm{CharMatch}(M(x),M^{\prime}(x))<\theta]$ be the true failure rate. Each verification trial is Bernoulli($p$). Observing $n$ successes (all pass), the likelihood ratio test gives: $\Pr[\text{all pass}\mid p]=(1-p)^{n}$. For $p\geq\epsilon$: $(1-\epsilon)^{n}\leq e^{-n\epsilon}$. Setting the false-acceptance probability to $e^{-n\epsilon}\leq e^{-2n\theta^{2}}$ (using $\epsilon=2\theta^{2}$ as the critical value from the Hoeffding bound on the CharMatch statistic), we obtain the stated bound. The Hoeffding refinement uses the fact that CharMatch is a bounded $[0,1]$ statistic with expected value $\geq\theta$ under $H_{0}$ (same model), applying $\Pr[\bar{X}-\mu\leq-t]\leq\exp(-2nt^{2})$ to the complement event. ∎∎
Theorem 10 (Credential Isolation).
@@
Proof.
Theorem 11 (Ledger Integrity).
@@
Proof.
We proceed by induction on chain position.
[…]

### 19 Mar 2026 v2

@@ 4.4 Reproducibility as a Trust Primitive
The threshold $\theta$ is task-type dependent: our empirical validation (Experiment 5, §7) shows that code-generating agents achieve perfect determinism ($\theta=0.98$), while prose-generating agents require $\theta=0.10$ to account for paraphrase-level variation. Cross-provider mean CharMatch ($0.114$, Table 5) remains well below even the relaxed threshold.
@@ 4.7 Security Analysis
Property 1 (Reproducibility Soundness, informal).
@@
By the DiFR token divergence result [1], the match rate between different models is $\delta\approx 0.06$ (character-level, validated in Experiment 5). Our empirical data (Table 6) shows that the fraction of cross-model output pairs exceeding any reasonable threshold $\theta$ is at most $q\approx 0.02$ — i.e., a substituted model’s output coincidentally passes the CharMatch $\geq\theta$ check with probability at most $q$ per prompt. Over $n$ independent verification prompts, the probability of all checks passing despite model substitution is at most $q^{n}$, which is negligible for realistic $n$.
Input integrity for replay verification. Replay verification requires the recorded input to be authentic. A compromised agent cannot fabricate its own input because the interaction ledger uses bilateral signing: the sender (an honest agent or non-agent) signs the input commitment $\eta_{\mathit{in}}$, which the compromised receiver cannot modify without invalidating the sender’s signature. G2 therefore relies on sender-authenticated input, not self-reported input—a unidirectional dependency on G3, not a circular one. The formal argument is given in Property 106).
@@ 4.8 Indistinguishability Under Sufficient Verification
$\epsilon\leq 1-\alpha^{1/n}$
where $\alpha$ is the significance level. For small $\epsilon$, this simplifies to $\epsilon\lessapprox\ln(1/\alpha)/n$. Equivalently, for target failure-rate bound $\epsilon$, the required verification budget is $n\geq\ln(1/\alpha)/\epsilon$.
Proof sketch. Let $p=\Pr_{x\sim\mathcal{D}}[\mathrm{CharMatch}(M(x),M^{\prime}(x))<\theta]$ be the true failure rate under $M^{\prime}$. Each verification trial $Z_{i}=\mathbf{1}[\mathrm{CharMatch}\geq\theta]$ is an independent Bernoulli(1p)(1-p). Observing all $n$ tests pass ($\sum Z_{i}=n$), the likelihood is $(1-p)^{n}$. The exact one-sided Clopper-Pearson upper bound on $p$ at confidence $1-\alpha$ is $p\leq 1-\alpha^{1/n}$. For practical computation: $(1-p)^{n}\leq\alpha$ implies $p\leq 1-\alpha^{1/n}\approx\ln(1/\alpha)/n$ when $p$ is small. This is a distribution-free result: the bound depends only on the number of verification trials and the desired confidence level, not on the threshold $\theta$ itself. The role of $\theta$ is to define what constitutes a “pass” — a higher $\theta$ increases sensitivity (more failures for a substituted model) but also increases the false-positive rate for the legitimate model. Full proof in §6. ∎
Practical implications. Table 2 shows the failure-rate bound $\epsilon$ (i.e., the maximum fraction of deployment prompts on which $M^{\prime}$ can diverge undetected) as a function of the verification budget $n$ at 99% confidence ($\alpha=0.01$):
$n$ $\epsilon$ bound Interpretation
10 0.369 Coarse screening
25 0.168 Moderate assurance
50 0.089 <<10% divergent prompts
100 0.045 <<5% divergent prompts
200 0.023 High assurance
500 0.009 <<1% divergent prompts
Table 2: Failure-rate bound $\epsilon$ at 99% confidence ($\alpha=0.01$) for various verification budgets $n$, computed as $\epsilon=1-0.01^{1/n}$.
A verification suite of 50 prompts per skill — operationally feasible as a periodic audit — bounds the adversary’s undetected divergence to at most 8.9% of deployment prompts with 99% confidence. Increasing to 200 prompts tightens this to 2.3%.
Threshold calibration. The threshold $\theta$ must be calibrated per task type and model to balance detection sensitivity against the false-positive rate. Our empirical data (Experiment 5) shows that single-metric thresholds suffer from the bimodal output distribution: for code generation, CharMatch is either >0.95>0.95 (identical outputs) or <0.5<0.5 (paraphrase variants), leaving no single $\theta$ that simultaneously achieves low false positives and high detection. We therefore adopt a multi-metric ensemble: a prompt is flagged if any of four metrics (CharMatch, Jaccard word overlap, TF-IDF cosine, character n-gram cosine) falls below its respective per-metric optimal threshold (determined by Youden’s $J$ statistic, Table 7 in §7). This ensemble approach improves detection robustness across task types, as each metric captures different aspects of output similarity (surface-form identity vs. semantic overlap).
Security interpretation. This transforms reproducibility from a detection mechanism into a safety certificate: an adversary who fine-tunes $M$ into $M^{\prime}$ that passes $n$ diverse replay checks has produced a model that diverges on at most an $\epsilon$-fraction of deployment prompts, with $\epsilon$ decreasing as $O(\ln(1/\alpha)/n)$. The adversary faces a dilemma: either $M^{\prime}$ behaves like $M$ on most inputs (limiting the attack’s impact to at most an $\epsilon$-fraction of requests), or $M^{\prime}$ diverges more broadly and will be caught with probability increasing in $n$.
@@ 4.9 Chain Verifiability
Per-model reproducibility variance. Our multi-provider experiment (Experiment 5c, §7) reveals that model determinism varies by over an order of magnitude: from CharMatch==0.987 (Haiku 4.5) to 0.169 (Kimi K2.5). Combined with the Chain Verifiability Theorem, this transforms into prescriptive architectural guidance: the security of a multi-agent pipeline depends not just on whether G2 is enforced, but on where high- and low-determinism models are placed in the trust tree.
Reproducibility as a model performance metric. The G1–G3 framework elevates reproducibility from an implementation detail to a first-class model performance metric: a model’s determinism directly determines which governance level it qualifies for (§8.2), which trust-tree positions it may occupy, and what verification budget is needed (Theorem 1). We argue that providers should self-declare reproducibility characteristics—analogous to FIPS 140 security levels for cryptographic modules—including expected CharMatch ranges under temperature==0 and sensitivity to infrastructure changes. Our empirical characterization across 9 models and 7 providers (Experiment 5) provides the first multi-provider basis for such specifications.
@@
Principle 1: High-determinism models at interior nodes. Models with strong reproducibility (Haiku: CharMatch==0.987, Sonnet: 0.823) occupy interior trust-tree positions; low-determinism models (Kimi: 0.169, DeepSeek: 0.225) are constrained to leaf positions—a security architecture constraint, not merely a performance recommendation.
@@
Principle 2: Verification boundary checkpoints. When a G2==none agent must appear at an interior position, the chain requires a verification boundary—human-in-the-loop, TEE attestation, or dual-execution—to re-establish chain verifiability for downstream nodes.
@@
Principle 3: Explicit degradation in the ledger. When chain verifiability breaks at node $v_{k}$, the ledger records a PARTIAL_VERIFIABILITY marker; downstream agents inherit a trust tier penalty, making the reduced guarantee visible to auditors.
Relationship to TEE. The Chain Verifiability Theorem clarifies where G2 is needed: at every interior node. In mixed deployments, different nodes may use different G2 implementations (TEE, software replay, or DV-SNARK), with chain verifiability depth determined by the weakest interior node. Regardless of the G2 implementation, each node still requires G1 and G3: TEE attests that some model ran correctly, but only G1 ensures it was the authorized model, and only G3 provides tamper-evident records.
4.10 Enhanced G2: Zero-Knowledge Behavioral Verification
The basic G2 instantiation requires provider cooperation: the verifier must re-execute the agent’s declared model on recorded inputs and compare outputs. This creates two limitations: (1) the verifier must have API access to the same model version, and (2) the original prompts and outputs are disclosed to the verifier, which may contain sensitive information.
Designated-Verifier SNARKs (DV-SNARKs) address both limitations by enabling the agent (or a delegated prover) to demonstrate output consistency in zero knowledge: the verifier learns only that $\mathrm{CharMatch}(O,O^{\mathrm{replay}})\geq\theta$, without seeing $O$, $O^{\mathrm{replay}}$, or the prompts.
Key design insight. Unlike BAID [2], which SNARKs over the entire transformer forward pass (${\sim}120$ s per proof for a 7B model), our approach SNARKs only over the verification metric computation—a fundamentally simpler circuit:
Definition 9 (CharMatch Verification Circuit).
$\mathcal{C}_{\mathrm{verify}}(O,O^{\mathrm{replay}},\theta)=\begin{cases}1&\text{if }\frac{|\{i:O[i]=O^{\mathrm{replay}}[i]\}|}{\max(|O|,|O^{\mathrm{replay}}|)}\geq\theta\\ 0&\text{otherwise}\end{cases}$
For typical agent outputs ($|O|\leq 4{,}096$ characters), the circuit requires $O(|O|)$ equality comparisons and a single division—on the order of $10^{4}$ R1CS constraints (specifically: 19,238 for 1K outputs, ${\sim}73{,}000$ for 4K outputs), compared to ${\sim}10^{10}$ for a transformer circuit. This reduces proof generation from minutes to sub-seconds.
Designated-Verifier property. The proof is bound to a specific verifier’s public key, preventing proof transfer. Only the designated auditor can verify the proof, which prevents a compromised agent from shopping for a lenient verifier. The DV property also prevents the proof from being used as transferable evidence—a deliberate design choice that limits the proof’s scope to the specific audit context.
Protocol. Enhanced G2 verification proceeds as:
  1. 1.
    The agent (or a trusted replay service) executes the model on the recorded input to produce $O^{\mathrm{replay}}$.
  2. 2.
    The prover computes $\pi=\mathrm{DV\text{-}SNARK}.\mathrm{Prove}(\mathcal{C}_{\mathrm{verify}},(O,O^{\mathrm{replay}}),\theta,\mathit{pk}_{\mathrm{verifier}})$.
  3. 3.
    The verifier checks $\mathrm{DV\text{-}SNARK}.\mathrm{Verify}(\pi,\theta,\mathit{sk}_{\mathrm{verifier}})\stackrel{{\scriptstyle?}}{{=}}1$.
What DV-SNARK does and does not prove. The SNARK proves that the prover possesses two strings whose CharMatch exceeds $\theta$; it does not prove that $O^{\mathrm{replay}}$ was produced by the declared model (that trust is delegated to the replay service). The SNARK’s contribution is privacy; for full computational integrity, TEE or BAID-style full-inference SNARKs remain necessary.
Cost-privacy tradeoff. DV-SNARK proof generation costs ${\sim}174$ ms (1K outputs) to ${\sim}750$ ms (4K outputs); Groth16 verification is constant at ${\sim}1.3$ ms (§7). Proof generation is ${\sim}1{,}740\times$ slower than basic CharMatch but is offline and asynchronous; it is ${\sim}690\times$ faster than BAID’s full-inference zkVM (${\sim}120$ s). The governance contract (G2) is identical; the privacy properties differ.
Security. Knowledge soundness ensures that a prover cannot produce a valid proof without possessing strings satisfying the CharMatch predicate. The DV property ensures simulation-privacy: no entity other than the designated verifier can distinguish a real proof from a simulated one.
@@ 5.1 Motivation
Why OpenTelemetry traces are insufficient. The industry standard for agentic observability is OpenTelemetry (OTEL) [47], whose span trees record execution traces with span_id/parent_id linkage, timestamped input.value/output.value payloads, and span-kind annotations (CHAIN, LLM, TOOL). Analysis of the Nemotron-AIQ Agentic Safety Dataset [46]—10,796 OTEL traces from NVIDIA’s AI-Q research assistant under 22 distinct attack templates—reveals three critical gaps: (1) No integrity protection: span linkage is purely referential; any party with storage access can modify, delete, or reorder spans undetectably. (2) No identity binding: spans record function names but not cryptographically verified agent identities or capability certificates. (3) No privacy preservation: OTEL stores full plaintext input/output payloads, creating data exposure risk. Our interaction ledger addresses all three gaps through hash-chain linking ($h_{\mathit{prev}}$), bilateral digital signing ($\mathit{sig}_{i}$; Ed25519 in the basic instantiation), agent certificate binding ($\mathit{ch}_{\mathit{send}},\mathit{ch}_{\mathit{recv}}$), and commitment-only storage ($\eta_{\mathit{in}},\eta_{\mathit{out}}$).
@@ 5.4 Forensic Reconstruction Protocol
5.5 Chain Auditability
Analogous to the Chain Verifiability Theorem for G2 (§4.9), interaction auditability is a chain property: a single missing interaction record along an execution path breaks forensic reconstruction for all downstream nodes on that path.
Definition 12 (Chain Auditability Depth).
[…]
with $\mathrm{CAD}(v_{n})=n$ if all records are present (fully auditable chain).
Property 2 (Chain Auditability).
If $\mathrm{CAD}(v_{n})=k<n$, then for all $i\geq k$:
  1. 1.
    Provenance gap: the auditor cannot trace $v_{i}$’s output back to the trust anchor $v_{0}$, because the input $v_{k}$ received from $v_{k-1}$ is unrecorded.
  2. 2.
    G2 cascade: replay verification for $v_{k}$ is impossible—the sender-authenticated input $\eta_{\mathit{in}}$ required by Property 10 is unavailable. By induction, downstream G2 verification inherits the same gap even if $\rho_{v_{i}}.\mathit{level}=\mathit{full}$.
This creates a symmetric pair of chain properties: G2’s Chain Verifiability (Theorem 2) governs behavioral verification depth, while G3’s Chain Auditability governs forensic reconstruction depth. Because G2 replay depends on G3’s sender-signed input (Property 10), the effective verification depth of any execution path is $\min(\mathrm{CVD},\mathrm{CAD})$—the minimum of the two chain depths.
5.6 G3 as a Governance Contract
Like G1 and G2, G3 defines what properties interaction records must satisfy, not how the ledger is implemented. The governance contract requires: (1) every inter-agent and agent-human interaction produces a record with bilateral signatures and content commitments; (2) records are linked in a tamper-evident chain; (3) sufficient metadata exists for forensic reconstruction.
Our current instantiation uses a centralized append-only ledger operated by a trusted organizational entity (e.g., the deploying enterprise). This is the simplest design that satisfies G3: agents submit bilaterally-signed records to the operator, who maintains the hash chain and provides query access during audits. The operator is assumed honest for availability (it will not selectively omit records) and ordering (it will not reorder records), while integrity is enforced cryptographically regardless of operator behavior (bilateral signatures + hash chain prevent undetectable modification).
Alternative instantiations satisfying the same G3 contract include: (a) multi-operator consensus, where f+1f{+}1-of-$n$ operators must agree on each append, eliminating single-operator availability trust at the cost of consensus latency; (b) append-only transparency logs analogous to Certificate Transparency, where multiple independent monitors cross-check the ledger for consistency; (c) blockchain-anchored commitments, where periodic ledger root hashes are anchored to a public blockchain for non-repudiable timestamping. Each alternative trades latency or operational complexity for stronger trust assumptions. The G3 contract is invariant across all instantiations—what changes is the trust model of the storage layer, not the governance semantics of the records.
6 Formal Security Properties
We state seven security properties and two theorems under the adversary model of §2. The basic instantiation uses standard cryptographic assumptions: EUF-CMA security of Ed25519, collision resistance of SHA-256, and the DiFR token divergence property [1]. The enhanced instantiation additionally relies on $q$-SDH and DDH in Type-3 bilinear groups (BBS+ unforgeability and zero-knowledge, §3.5) and knowledge soundness of Groth16 (DV-SNARK, §4.10). All proofs are in Appendix A.
Property 3 (Trust Containment).
@@ 6 Formal Security Properties
Property 4 (Capability Binding).
@@
Property 5 (Chain Integrity).
@@
Property 6 (Delegation Depth Enforcement).
@@
Property 7 (Reproducibility Soundness).
Under software G2 (replay verification): if $v$ executes $M^{\prime}\neq M$ or $S^{\prime}\neq S$, replay verification detects with probability $\geq 1-\epsilon(\lambda)$. Under hardware G2 (TEE): detection is deterministic.
@@
If $M^{\prime}$ passes replay verification (CharMatch $\geq\theta$) on $n$ independent prompts from $\mathcal{D}$, then $\Pr_{x\sim\mathcal{D}}[\mathrm{CharMatch}(M(x),M^{\prime}(x))<\theta]\leq\epsilon$ where $\epsilon\leq 1-\alpha^{1/n}$, with $\alpha$ the significance level.
Property 8 (Credential Isolation).
@@
Property 9 (Ledger Integrity).
@@
Property 10 (G2 Input Integrity via Bilateral Signing).
Replay verification for a compromised agent $v_{j}$ uses sender-authenticated input, not self-reported input. Consequently, G2 soundness does not circularly depend on $v_{j}$’s own honesty.
Rationale. A potential concern is circular dependency: G2 replay verification requires the recorded input to be authentic, but if the compromised agent $v_{j}$ controls its own input recording, it could fabricate inputs that make its unauthorized output appear legitimate. We show this is precluded by the bilateral signing structure of G3.
In any execution chain, agent $v_{j}$’s input is the output of its predecessor $v_{j-1}$. The interaction record $R_{i}$ in the ledger contains the input commitment $\eta_{\mathit{in}}=\mathrm{SHA\text{-}256}(\mathit{input})$ and is bilaterally signed by both $v_{j-1}$ (sender) and $v_{j}$ (receiver). We consider three cases for $v_{j-1}$:
Case 1: $v_{j-1}$ is a non-agent node (human or organization). By the adversary model (§2), non-agent nodes are uncompromised. $v_{j-1}$ signs the interaction record containing $\eta_{\mathit{in}}$. Agent $v_{j}$ cannot modify $\eta_{\mathit{in}}$ without invalidating $v_{j-1}$’s signature (EUF-CMA). Therefore the input committed to the ledger is the true input $v_{j}$ received.
Case 2: $v_{j-1}$ is an honest agent. The same argument applies: $v_{j-1}$ signs its output and the corresponding ledger record. $v_{j}$ cannot forge $v_{j-1}$’s signature on a different $\eta_{\mathit{in}}$.
Case 3: $v_{j-1}$ is also compromised. Then $v_{j-1}$’s output is itself unverifiable. However, the Chain Verifiability Theorem (Theorem 4) already establishes that all nodes downstream of the first unverifiable agent lose end-to-end verification guarantees. In this case, $v_{j}$ is downstream of the already-compromised $v_{j-1}$, so the chain verifiability depth $\mathrm{CVD}\leq j-1<j$, and no claim of end-to-end verification extends to $v_{j}$ regardless.
In Cases 1 and 2, the verifier reconstructs $v_{j}$’s input via the forensic reconstruction protocol (§5): request $v_{j-1}$ to disclose the original content, verify it against the sender-signed $\eta_{\mathit{in}}$, then replay $v_{j}$’s declared model on this authenticated input. The dependency is therefore unidirectional—G2 relies on G3’s sender-side input commitment, which is established independently of $v_{j}$—not circular.
Chain Auditability. This unidirectional dependency implies that G3 also exhibits a chain property: if an interaction record along an execution path is absent from the ledger, G2 replay verification at that point becomes impossible (the authenticated input is unavailable), and by induction all downstream provenance tracing fails. The effective verification depth of any execution path is therefore $\min(\mathrm{CVD},\mathrm{CAD})$—the minimum of Chain Verifiability Depth (Definition 8) and Chain Auditability Depth (Definition 12). This symmetric pair of chain properties—G2 governing behavioral verification depth, G3 governing forensic reconstruction depth—underscores the mutual reinforcement of the governance requirements.